safyousuff:

If 27.0 mL of Ca(OH)2 with an unknown concentration is neutralized by 32.40 mL of 0.185 M HCl, what is the concentration of the Ca(OH)2 solution? Show all of the work needed to solve this problem. Ca(OH)2 + 2 HCl yields 2 H2 O + CaCl2

JustSaiyan:

Ca(OH)2(aq) + 2HCl(aq) → CaCl2(aq) + 2H2O(l) n(Ca(OH)2) = 1/2 · n(HCl) = (1/2)(0.03240 L)(0.185 M) = 0.003 mol [Ca(OH)2] = 0.003 mol ÷ 0.0270 L = 0.111 M 0.111 M

safyousuff:

thanks :)

JustSaiyan:

You're welcome.